OpenStudy (bigbosssaint21):

Choose the polynomial that is written in standard form. 2x^3y^2 + 4x^4y + 8x^7 -3x^4y^2 + 4x^4y^5 + 10x^2 6xy^2 + 4x^3y^8 + 9x^2 -2x^5y^3 + 4x^2y^2 + x^2

OpenStudy (bigbosssaint21):

@vzfreakz

OpenStudy (anonymous):

Lo, you have another question?

OpenStudy (anonymous):

lol*

OpenStudy (bigbosssaint21):

yea lol i got several

OpenStudy (bigbosssaint21):

this is a new one

OpenStudy (anonymous):

YES! I got my Human Calculator title :)

OpenStudy (bigbosssaint21):

lol

OpenStudy (anonymous):

Ih dang it! I knew I should have not bought that yet. Trying to get Mathlete.

OpenStudy (anonymous):

I would go with A. But not positive because as I know that finding the standard form you have to put the terms with the highest degree first.

OpenStudy (anonymous):

A or C.

OpenStudy (bigbosssaint21):

ok ready for another?

OpenStudy (bigbosssaint21):

Identify the Exponent Law shown in the example below. a2a3 = a2+3 = a5 Power of a Product Power of a Power Product of Powers None of the above

OpenStudy (bigbosssaint21):

a^2a^3=a^2+3=a^5

OpenStudy (anonymous):

Product of power.

OpenStudy (bigbosssaint21):

Choose the correct simplification of the expression (2ab^5)^3. 8ab^15 8a^3b^8 5a^3b^15 8a^3b^15

OpenStudy (anonymous):

Let's keep track on the medals now.

OpenStudy (anonymous):

First one.